OpenStudy (anonymous):

What is the sum of 7/x-1+9/x-8 ?

OpenStudy (unklerhaukus):

\[\frac7{x-1}+\frac9{x-8}\\=\frac7{x-1}\times\frac{x-8}{x-8}+\frac9{x-8}\times\frac{x-1}{x-1}\\=\frac{7(x-8)+9(x-1)}{(x-1)(x-8)}\\=\\=\]